OpenStudy (anonymous):

I know there's a formula for this, I just have no idea what it is :( Find the missing term of the following sequence. ...-2,_,456,... A.227 B.221 C.219 D.243

OpenStudy (anonymous):

-2+2d=456 2d=456+2=458 d=458/2=229 miising term = -2+d=-2+229=227 answer A.

OpenStudy (anonymous):

Oh... What i was doing was just adding each answer to the -2 twice. Thanks a lot!

OpenStudy (anonymous):

yw